### v1

@@ Abstract
In this work, we propose a noise-centric formulation of the distributionally robust Kalman filter (DRKF) for discrete-time linear stochastic systems with uncertain noise statistics. By placing Wasserstein ambiguity sets directly on the process and measurement noise distributions, the proposed DRKF preserves the analytical structure of the classical Kalman filter while providing a priori spectral bounds on all feasible covariances. In the time-invariant setting, we derive a steady-state DRKF from a single stationary semidefinite program, yielding a constant-gain estimator with the same per-step computational complexity as the standard Kalman filter. We establish conditions guaranteeing the existence, uniqueness, and convergence of this steady-state solution, and we prove its asymptotic minimax optimality with respect to the worst-case mean-square error. Numerical experiments validate the theory and demonstrate that the proposed DRKF improves estimation accuracy under unknown or uncertain noise models while offering computational advantages over existing robust and distributionally robust filters.
@@ 1 Introduction
State estimation is a central problem in control, perception, signal processing, and robotics, where the goal is to infer system states from noisy and partial measurements. The Kalman filter (KF) [1] is optimal for linear–Gaussian systems when the noise statistics are known. In practice, however, these statistics are rarely available a priori; they are estimated from limited data, subject to modeling error, and often vary over time. Even mild misspecification can degrade estimation accuracy, while severe mismatch may lead to filter divergence.
@@
To address this sensitivity, several robust estimation frameworks have been proposed. The $H_{\infty}$ filter [2] bounds the disturbance-to-error energy gain and guarantees finite error under worst-case noise, but it is often conservative in stochastic settings due to its pessimistic nature. Risk-sensitive filters [3, 4] penalize large errors exponentially, but their performance depends on a risk parameter that lacks a clear statistical interpretation and is difficult to tune, frequently resulting in excessive conservatism or instability.
Recent advances in distributionally robust optimization (DRO) have inspired the development of distributionally robust state estimators (DRSE). Although distributional robustness has already evolved into a rich research area in control design (e.g., [5, 6, 7, 8, 9, 10, 11]), its exploration in state estimation remains relatively limited. DRSE methods generally assume that the true noise distributions lie within an ambiguity set centered at a nominal distribution and seek estimators that minimize the worst-case expected error over this set. Within this framework, a variety of ambiguity sets have been considered, including $\phi$-divergence balls, moment-based ambiguity sets [12, 13, 14, 15], and Wasserstein balls [16, 17, 18, 19, 20]. Importantly, DRSE formulations differ not only in the type of ambiguity set employed, but also in the distributions to which the ambiguity is applied. For example, [16] imposes a Wasserstein ambiguity set on the joint state–measurement distribution; [17] places separate ambiguity sets on the prior state and measurement-noise covariances; and [19] introduces a bicausal optimal-transport ambiguity set that explicitly encodes temporal causality.
@@
Despite this progress, steady-state formulations of DRSE remain underdeveloped. Analyzing the infinite-horizon regime is crucial for long-term operation and computational tractability: time-varying DRKFs typically require solving an optimization problem at every time step or a large optimization problem whose size grows with the horizon. The frequency-domain minimax formulation of [20] characterizes the optimal linear time-invariant estimator, but the resulting filter is non-rational and does not admit a finite-dimensional state-space realization.
@@
Our previous work [21] developed a steady-state DRKF that placed ambiguity sets on the prior state and measurement-noise distributions. In this paper, we instead adopt a noise-centric formulation, placing Wasserstein ambiguity sets directly on the initial state, process noise, and measurement noise distributions. Unlike existing approaches (e.g., [20]), the resulting steady-state DRKF is obtained from a single stationary semidefinite program (SDP) that directly returns a constant-gain solution. This noise-centric formulation ensures a form of dynamic consistency—all admissible priors are generated by propagating the system dynamics under feasible noise sequences—and enables several theoretical and computational advantages.
@@
  • Noise-centric ambiguity modeling. Wasserstein ambiguity sets are defined directly on the initial state, process noise, and measurement noise distributions. This design $(i)$ enforces a form of dynamic consistency by ensuring that all admissible priors are reachable through the system dynamics under admissible noise realizations; $(ii)$ provides a priori spectral bounds on the DRKF covariance recursion, unlike ambiguity sets using joint state–measurement laws, which may include unreachable priors and typically lack such spectral guarantees; and $(iii)$ yields a tractable time-varying DRKF whose stage-wise gains are precomputed via small convex SDPs and applied online with Kalman-level per-step complexity.
  • Steady-state DRKF via a single SDP. A constant-gain DRKF is obtained by solving one stationary SDP offline. Its online computational cost matches that of the classical KF. The resulting steady-state covariances solve a time-invariant convex program, and the gains of the time-varying DRKF converge to this stationary gain.
  • Theoretical guarantees. $(i)$ Spectral boundedness and KF sandwiching: Wasserstein balls impose explicit eigenvalue bounds on feasible covariances, ensuring that the DRKF’s prior and posterior covariances lie between those of two appropriately chosen KFs. This property prevents arbitrarily inflated or singular covariances under worst-case deviations in the noise statistics. $(ii)$ Explicit convergence radii: we derive closed-form sufficient conditions on the ambiguity radii guaranteeing contraction of the DR Riccati recursion and convergence of the time-varying DRKF to its steady-state counterpart. $(iii)$ Stability and optimality: under standard controllability and observability assumptions, the estimation error dynamics are Schur stable, and the steady-state DRKF achieves asymptotic minimax mean-square error (MSE) optimality within the specified ambiguity sets.
The remainder of this paper is organized as follows. Section˜2 reviews preliminaries on DRSE and Wasserstein ambiguity sets. Section˜3 presents the time-varying and steady-state DRKF algorithms and establishes their key properties. Section˜4 develops convergence, stability, and optimality results. Section˜5 demonstrates the performance and utility of the proposed framework through numerical experiments.
@@ 2 Problem Setup
We begin by introducing notation. Let $\mathbb{S}^{n}$ denote the set of symmetric matrices in $\real{n\times n}$, with $\mathbb{S}^{n}_{+}$ and $\mathbb{S}^{n}_{++}$ denoting the positive semidefinite and positive definite cones. For $A,B\in\mathbb{S}^{n}$, the relations $A\succeq B$ and $A\succ B$ indicate that $A-B\in\mathbb{S}^{n}_{+}$ and $A-B\in\mathbb{S}^{n}_{++}$, respectively. For $A\in\real{m\times n}$, $\|A\|_{2}$ and $\|A\|_{F}$ denote the spectral and Frobenius norms. For $A\in\mathbb{S}^{n}$, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote its smallest and largest eigenvalues, and $A^{+}$ denotes its Moore–Penrose pseudoinverse.
Let $\mathcal{P}(\mathcal{W})$ be the set of Borel probability measures supported on a set $\mathcal{W}\subseteq\real{n}$, and let $\mathcal{P}_{2}(\mathcal{W})$ denote the subset of measures with finite second-order moments. Given two probability measures $\mathbb{P}$ and $\mathbb{Q}$, their product measure is denoted by $\mathbb{P}\otimes\mathbb{Q}$. For a sequence $\{w_{k}\}_{k\geq 0}$ with $w_{k}\in\real{n}$, the stacked block vector of length $N$ at time $t$ is defined as $\bm{w}_{t}^{N}\coloneqq[\,w_{tN+N-1}^{\top},\dots,w_{tN}^{\top}\,]^{\top}$.
@@ 2.1 Distributionally Robust State Estimation Problem
Consider the discrete-time linear stochastic system [^1] Throughout, we assume that the system matrices $A$ and $C$ are known, and we focus exclusively on uncertainty in the noise statistics.
@@
where $x_{t}\in\real{n_{x}}$ and $y_{t}\in\real{n_{y}}$ denote the system state and the output. The process noise $w_{t}\in\real{n_{x}}$ and the measurement noise $v_{t}\in\real{n_{y}}$ follow distributions $\mathbb{Q}_{w,t}\in\mathcal{P}(\real{n_{x}})$ and $\mathbb{Q}_{v,t}\in\mathcal{P}(\real{n_{y}})$, while the initial state $x_{0}$ follows $\mathbb{Q}_{x,0}\in\mathcal{P}(\real{n_{x}})$. We assume that $x_{0}$, $\{w_{t}\}$, and $\{v_{t}\}$ are mutually independent, and that each of $\{w_{t}\}$ and $\{v_{t}\}$ is temporally independent; that is, $w_{t}$ and $w_{t^{\prime}}$ (and likewise $v_{t}$ and $v_{t^{\prime}}$) are independent for all $t\neq t^{\prime}$. [^2] Such independence assumptions are standard in the control and estimation literature; see, e.g., [22, Sec. 2.2] and [2, Sec. 5.1].
At each time $t\geq 0$, let $\mathcal{Y}_{t}\coloneqq\{y_{0},\dots,y_{t}\}$ denote the measurement history. The goal is to estimate $x_{t}$ given $\mathcal{Y}_{t}$. Since outputs arrive sequentially, a natural approach is to compute the minimum mean-square error (MMSE) estimator recursively:
@@
where $\mathcal{F}_{t}=\{\,\psi_{t}:(\real{n_{y}})^{t+1}\rightarrow\real{n_{x}}:\psi_{t}\text{ measurable}\,\}$ denotes the set of admissible estimators. The expectation is taken with respect to the conditional joint distribution of $(x_{t},y_{t})$ given the past observations $\mathcal{Y}_{t-1}$. Under (1), this conditional distribution is fully determined, for $t>0$, by the posterior distribution $\mathbb{P}_{x,t-1}$ of $x_{t-1}$ given $\mathcal{Y}_{t-1}$ and the joint noise distribution $\mathbb{Q}_{e,t}\coloneqq\mathbb{Q}_{w,t-1}\otimes\mathbb{Q}_{v,t}$ of $(w_{t-1},v_{t})$. At $t=0$, no posterior from a previous step exists. By slight abuse of notation, we adopt the convention $\mathcal{Y}_{-1}=\emptyset$ and set $\mathbb{P}_{x,0}^{-}\coloneqq\mathbb{Q}_{x,0}$ and $\mathbb{Q}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{Q}_{v,0}$, so that the definition in (2) also covers $t=0$, where the expectation reduces to an unconditional expectation with respect to the joint law of $(x_{0},v_{0})$. [^3] We use the superscript “-” for prior quantities; e.g., $\bar{x}^{-}_{t}$, $\Sigma_{x,t}^{-}$, and $\mathbb{P}_{x,t}^{-}$ denote the prior mean, covariance, and distribution of $x_{t}$ conditioned on $\mathcal{Y}_{t-1}$.
When $x_{0}\sim\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,t}^{-})$ and $\{w_{t}\},\{v_{t}\}$ are independent Gaussian sequences with known covariances, the optimal causal MMSE estimator reduces to the classical KF. In practice, noise statistics are estimated from limited data and thus subject to ambiguity. Typically only nominal distributions $\hat{\mathbb{Q}}_{w,t}$, $\hat{\mathbb{Q}}_{v,t}$, and $\hat{\mathbb{Q}}_{x,0}$ are available from identification methods [23, 24, 25, 26]. [^4] A hat, $\hat{\cdot}$, denotes nominal quantities; e.g., $\hat{x}_{0}^{-}$ and $\hat{\mathbb{Q}}_{x,0}$ represent the nominal initial state mean and distribution, respectively. Even modest deviations between nominal and true distributions can degrade performance or cause divergence, especially when disturbances accumulate over time (e.g., marginally stable or unstable systems).
To address this, we formulate a stage-wise minimax problem in which the estimator competes against an adversary that selects distributions from ambiguity sets $\mathbb{D}_{j,t}\subset\mathcal{P}(\real{n_{j}})$ for $j\in\{w,v,x_{0}\}$. Each ambiguity set is centered at its corresponding nominal distribution and parameterized by a robustness radius. The DRSE problem at time $t$ is then given by
@@
where $\mathbb{D}_{e,0}\coloneqq\mathbb{D}_{x,0}\times\mathbb{D}_{v,0}$ and $\mathbb{D}_{e,t}\coloneqq\mathbb{D}_{w,t-1}\times\mathbb{D}_{v,t}$ for $t>0$. The corresponding joint distributions are given by $\mathbb{P}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{P}_{v,0}$ and $\mathbb{P}_{e,t}\coloneqq\mathbb{P}_{w,t-1}\otimes\mathbb{P}_{v,t}$. This product form reflects the assumed independence between process and measurement noise at each stage.
@@
In contrast to [17] and [21], which define the ambiguity set around the joint prior-measurement distribution, our formulation places ambiguity directly on the noise distributions. This enforces a form of dynamic consistency: priors arise only through the system dynamics (1) and admissible noise realizations. By contrast, DR-MMSE formulations may allow priors that are not dynamically reachable. Our formulation also enables the adversary to couple estimation performance across time, capturing long-term accumulation of disturbances via the state recursion, even though the ambiguity sets themselves are stage-wise. For systems with slow or unstable dynamics, even small biases in $w_{t}$ can lead to unbounded state growth, making this sequential robustness essential.
While DRSE formulations have been explored in several contexts [17, 18, 16], their application to control-oriented state estimation remains limited. Moreover, understanding their steady-state behavior as $t\to\infty$ is critical for assessing long-term robustness and computational tractability.

### v3

@@ Abstract
We study state estimation for discrete-time linear stochastic systems under distributional ambiguity in the initial state, process noise, and measurement noise. We propose a noise-centric distributionally robust Kalman filter (DRKF) based on Wasserstein ambiguity sets imposed directly on these distributions. This formulation excludes dynamically unreachable priors and yields a Kalman-type recursion driven by least-favorable covariances computed via semidefinite programs (SDP). In the time-invariant case, the steady-state DRKF is obtained from a single stationary SDP, producing a constant gain with Kalman-level online complexity. We establish the convergence of the DR Riccati covariance iteration to the stationary SDP solution, together with an explicit sufficient condition for a prescribed convergence rate. We further show that the proposed noise-centric model induces a priori spectral bounds on all feasible covariances and a Kalman filter sandwiching property for the DRKF covariances. Finally, we prove that the steady-state error dynamics are Schur stable, and the steady-state DRKF is asymptotically minimax optimal with respect to worst-case mean-square error.
@@ 1 Introduction
State estimation is a fundamental problem in systems and control, where the objective is to infer the system state from noisy and partial measurements. For linear systems with Gaussian noise and known statistics, the Kalman filter (KF) [12] provides the optimal minimum mean-square error (MMSE) estimator. In practice, however, noise statistics are rarely known a priori: they are typically learned from limited data, subject to modeling error, and may vary over time. Even moderate misspecification can significantly degrade estimation performance, while severe mismatch may lead to divergence.
@@
To mitigate this sensitivity, several robust estimation frameworks have been developed. The $H_{\infty}$ filter [27] guarantees bounded estimation error under worst-case disturbances, but it is often conservative in stochastic settings. Risk-sensitive filters (e.g., [28] and [32]) penalize large errors exponentially via an entropic risk measure, but their performance relies on an accurate specification of the underlying noise distributions, which may be difficult to justify in practice.
In contrast, recent advances in distributionally robust optimization (DRO) have motivated growing interest in distributionally robust state estimation (DRSE). In DRSE, the true noise distributions are assumed to lie in an ambiguity set centered at a nominal model, and the estimator minimizes the worst-case expected error over this set. While distributional robustness has been extensively studied in control design (e.g., [29, 33, 25, 17, 7, 19, 3]), its application to state estimation remains comparatively limited.
Existing DRSE formulations vary along two key dimensions: (i) the geometry of the ambiguity set, such as $\phi$-divergence balls, moment-based sets [30, 31, 35, 15], or Wasserstein balls [26, 23, 18, 8, 13]; and (ii) the location where ambiguity is imposed, e.g., on the joint state–measurement distribution [26], on the prior state and measurement noise [23], or via bicausal optimal transport models that encode temporal dependence [8]. These choices have significant implications for tractability, interpretability, and long-term behavior.
@@
Despite this progress, steady-state DRSE remains underdeveloped. For long-horizon operation, constant-gain estimators with fixed offline complexity are essential. However, many existing time-varying DRKFs require solving an optimization problem at each time step, or lead to horizon-dependent formulations whose size grows with time. While the frequency-domain minimax approach of [13] characterizes an optimal linear time-invariant estimator, the resulting filter is generally non-rational and does not admit a finite-dimensional state-space realization.
@@
The preliminary version [11] of this work derived a steady-state DRKF by placing ambiguity sets on the prior state and measurement noise distributions. In this paper, we substantially extend this line of work by adopting a noise-centric formulation, in which Wasserstein ambiguity sets are imposed directly on the initial state, process noise, and measurement noise distributions. This modeling choice enforces a form of dynamic consistency: all admissible prior distributions arise through propagation of the system dynamics under admissible noise realizations. Moreover, it admits a finite-dimensional steady-state characterization and yields a constant-gain DRKF from a single stationary semidefinite program (SDP), while providing several theoretical and computational advantages.
@@
  • Noise-centric ambiguity modeling. We define Wasserstein ambiguity sets directly on the initial state, process noise, and measurement noise distributions. This modeling choice enforces dynamic consistency, ensuring that all admissible prior distributions arise through propagation of the system dynamics under admissible noise realizations, and forms the structural basis for the subsequent theoretical and algorithmic developments.
  • Steady-state DRKF via a single stationary SDP. In the time-invariant setting, we obtain a constant-gain steady-state DRKF by solving a single stationary SDP offline, yielding Kalman-level per-step online complexity.
  • Convergence with explicit rate conditions. We establish the convergence of the DR Riccati iteration governing the state-error covariance to a unique steady-state solution characterized by the stationary SDP. Moreover, we derive an explicit sufficient condition on the ambiguity radii that guarantees a prescribed contraction rate.
  • Spectral structure and steady-state guarantees. We show that Wasserstein ambiguity induces explicit a priori eigenvalue bounds (an eigenvalue tube) and a KF-sandwich property for the DRKF covariances. We further establish Schur stability of the steady-state estimation error dynamics and asymptotic minimax optimality with respect to the worst-case mean-square error (MSE).
The remainder of the paper is organized as follows. Section˜2 introduces the problem formulation and Wasserstein ambiguity sets. Section˜3 presents the finite-horizon and steady-state DRKF constructions and their SDP reformulations. Section˜4 develops spectral boundedness, convergence, stability, and optimality results. Section˜5 illustrates the performance of the proposed methods through numerical experiments.
@@ 2 Problem Setup
Notation. Let $\mathbb{S}^{n}$ denote the set of symmetric matrices in $\real{n\times n}$, and let $\mathbb{S}^{n}_{+}$ and $\mathbb{S}^{n}_{++}$ be the positive semidefinite and positive definite cones. For $A,B\in\mathbb{S}^{n}$, write $A\succeq B$ (resp. $A\succ B$) if $A-B\in\mathbb{S}^{n}_{+}$ (resp. $\mathbb{S}^{n}_{++}$). For a matrix AA, $\|A\|_{2}$ and $\|A\|_{F}$ denote the spectral and Frobenius norms. For $A\in\mathbb{S}^{n}$, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote its smallest and largest eigenvalues, while $\sigma_{\min}(A)$ and $\sigma_{\max}(A)$ denote its smallest and largest singular values. Let $\mathcal{P}(\mathcal{W})$ be the set of Borel probability measures supported on $\mathcal{W}\subseteq\real{n}$, and let $\mathcal{P}_{2}(\mathcal{W})$ be those with finite second moments. For probability measures $\mathbb{P},\mathbb{Q}$, their product measure is denoted by $\mathbb{P}\otimes\mathbb{Q}$. Let $\mathbf{1}_{N}\in\mathbb{R}^{N}$ denote the vector of all ones, and let $\operatorname{blkdiag}(A_{1},\dots,A_{N})$ denote the block-diagonal matrix with diagonal blocks $A_{1},\dots,A_{N}$.
@@ 2.1 Distributionally Robust State Estimation Problem
Consider the discrete-time linear stochastic system [^1] Throughout, we assume that the system matrices $A_{t}$ and $C_{t}$ are known, and we focus exclusively on uncertainty in the noise statistics.
@@
where $x_{t}\in\real{n_{x}}$ and $y_{t}\in\real{n_{y}}$ denote the system state and output. The process noise $w_{t}\in\real{n_{x}}$ and measurement noise $v_{t}\in\real{n_{y}}$ follow distributions $\mathbb{Q}_{w,t}\in\mathcal{P}(\real{n_{x}})$ and $\mathbb{Q}_{v,t}\in\mathcal{P}(\real{n_{y}})$, while the initial state $x_{0}$ follows $\mathbb{Q}_{x,0}\in\mathcal{P}(\real{n_{x}})$. We assume that $x_{0}$, $\{w_{t}\}$, and $\{v_{t}\}$ are mutually independent, and that each of $\{w_{t}\}$ and $\{v_{t}\}$ is temporally independent. [^2] Such independence assumptions are standard in the control and estimation literature; see, e.g., [1, Sect. 2.2] and [27, Sect. 5.1].
At each time $t\geq 0$, the estimator has access to the measurement history $\mathcal{Y}_{t}\coloneqq\{y_{0},\dots,y_{t}\}$. The goal is to estimate $x_{t}$ given $\mathcal{Y}_{t}$. A natural criterion is the conditional minimum mean-square error (MMSE),
@@
where $\mathcal{F}_{t}$ denotes the set of measurable estimators $\psi_{t}:(\real{n_{y}})^{t+1}\to\real{n_{x}}$. That is, $\psi_{t}$ minimizes the conditional mean-square estimation error given past observations.
For $t>0$, the conditional distribution of $(x_{t},y_{t})$ given $\mathcal{Y}_{t-1}$ is fully determined by the posterior distribution $\mathbb{P}_{x,t-1}$ of $x_{t-1}$ and the joint noise distribution $\mathbb{Q}_{e,t}\coloneqq\mathbb{Q}_{w,t-1}\otimes\mathbb{Q}_{v,t}$. At $t=0$, no posterior from a previous step exists. We adopt the convention $\mathcal{Y}_{-1}=\emptyset$ and define $\mathbb{P}_{x,0}^{-}\coloneqq\mathbb{Q}_{x,0}$ and $\mathbb{Q}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{Q}_{v,0}$, so that (2) also covers the initial stage. [^3] We use the superscript “-” for prior quantities; e.g., $\bar{x}^{-}_{t}$, $\Sigma_{x,t}^{-}$, and $\mathbb{P}_{x,t}^{-}$ denote the prior mean, covariance, and distribution of $x_{t}$ conditioned on $\mathcal{Y}_{t-1}$.
When $x_{0}\sim\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,0}^{-})$ and $\{w_{t}\},\{v_{t}\}$ are independent Gaussian sequences with known covariances, the optimal causal MMSE estimator reduces to the classical KF. In practice, however, noise statistics are typically estimated from limited data and are thus subject to ambiguity. In practice, only nominal distributions $\hat{\mathbb{Q}}_{w,t}$, $\hat{\mathbb{Q}}_{v,t}$, and $\hat{\mathbb{Q}}_{x,0}$ are available from identification procedures (e.g., [20, 24]). [^4] A hat, $\hat{\cdot}$, denotes nominal quantities; e.g., $\hat{x}_{0}^{-}$ and $\hat{\mathbb{Q}}_{x,0}$ represent the nominal initial state mean and distribution, respectively. Even modest deviations between nominal and true distributions can significantly degrade performance or cause divergence, especially when disturbances accumulate over time (e.g., marginally stable or unstable systems).
To address this, we formulate a stage-wise minimax problem in which the estimator competes against an adversary that selects distributions from ambiguity sets $\mathbb{D}_{j,t}\subset\mathcal{P}(\real{n_{j}})$ for $j\in\{w,v,x\}$ to be defined later. The resulting DRSE problem at time $t$ is
@@
where $\mathbb{D}_{e,0}\coloneqq\mathbb{D}_{x,0}\times\mathbb{D}_{v,0}$ and $\mathbb{D}_{e,t}\coloneqq\mathbb{D}_{w,t-1}\times\mathbb{D}_{v,t}$ for $t>0$. The corresponding joint distributions are $\mathbb{P}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{P}_{v,0}$ and $\mathbb{P}_{e,t}\coloneqq\mathbb{P}_{w,t-1}\otimes\mathbb{P}_{v,t}$, reflecting the assumed stage-wise independence between process and measurement noise.
@@
In contrast to formulations that place ambiguity on a joint prior state–measurement distribution (e.g., [23, 11]), our approach imposes ambiguity directly on the noise distributions. This enforces a form of dynamic consistency: admissible priors arise only through the system dynamics (1) under admissible noise realizations. By contrast, previous DRSE formulations may allow priors that are not dynamically reachable. Our formulation also allows the adversary to couple estimation performance across time, capturing long-term accumulation of disturbances via the state recursion, even though the ambiguity sets themselves are stage-wise. For systems with slow or unstable dynamics, even small biases in $w_{t}$ can lead to unbounded state growth, making this sequential robustness essential.